what is the percent yield if you start with 25g C3H6 are produced when 33.5 g are used ?

C3H6 +Br2 >>C3H6Br2
Answer #1

formula

% yield = actual * 100 / theoretical mass

reaction

C3H6 + Br2 -------> C3H6Br2

molar mass of C3H6 MM = 3 * MM(C) + 6 * MM(H)

= 3 * 12.01 + 6* 1.008

= 42.078 g/mol

mass (C3H6) = 25g

number of moles of (C3H6) n = mass of (C3H6) / molar mass of(C3H6)

n = 25g / 42.078 g/mol

n = 0.594 mol

molar mass of C3H6Br2

MM = 3 * MM(C) + 6 * MM(H) + 2* MM(Br)

= 3 * 12.01 + 6* 1.008 + 2* 79.904

= 201.89 g/mol

according to the balanced equation

C3H6 + Br2 -------> C3H6Br2

mol of C3H6Br2 formed = moles of C3H6

= 0.594 mol

use mass of CO2 = number of mol* molar mass

= 0.594 * 201.89

= 119.92 g

% yield = actual * 100 / theoretical mass

= 33.5 * 100 / 119.92

= 27.9%

% yield = 27.9%

i hope this help you

give thums up if you like

thank you